will a tutor please help me,
Lin is now 10 years old. Her dad is 3 times as old In how many years will her dad be twice as old as Lin?How old will each be then?
years later-
Dad then-
Lin then-

Answer by Paul(988)

Lin is 10 years old so her dad is 30 years old.
Let x be the age in the futre.
Lin's age = 10+x
Her dad's age = x+30

Equation:

2(x+10)=x+30
2x+20=x+30
x=10

Hence, after 10 years.
Paul.